Firm A has a margin of 11%, sales of $570,000, and ROI of 19%. Calculate the...

  1. Firm A has a margin of 11%, sales of $570,000, and ROI of 19%. Calculate the firm's average total assets.
  2. Firm C has net income of $136,000, turnover of 2.01, and ROI of 23.70%. Calculate the firm's margin, sales, and average total assets.
Answer #1

Answer:

a) Turnover = 19/11 = 1.7273

Average total assets = Sales/Turnover

                                = 570000/1.7273

Average total assets = $330,000

b) Margin = 23.70/2.01 = 11.79%

Sales = 136000/.1179 = $11,53,520(approx)

Average total assets = 11,53,520/2.01 = $5,73,890.547

Note : please consider rounding of difference.